# fastkqr: A Fast Algorithm for Kernel Quantile Regression

### v1

@@ 3.1 Methodology
When KQR is fitted at multiple quantile levels individually, say, for $10<\tau_{1}<\tau_{2}<\cdots<\tau_{T}<$ 1,
@@
the fitted curves may cross each other. To avoid the occurrence of crossing, Takeuchi et al. (2006), Bondell et al. (2010), and Liu and Wu (2011) consider fitting all the quantile curves simultaneously with a hard non-crossing constraint,
@@
where the notation $\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T}$ represents the collection of $b_{\tau_{1}},\boldsymbol{\alpha}_{\tau_{1}},b_{\tau_{2}},\boldsymbol{\alpha}_{\tau_{2}},\cdots,b_{\tau_{T}},\boldsymbol{\alpha}_{\tau_{T}}$ for ease of presentation. In the sequel, we extend this notation to $\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}_{t=1}^{T}$ and $\{\tilde{b}_{\tau_{t}},\tilde{\boldsymbol{\alpha}}_{\tau_{t}}\}_{t=1}^{T}$ to represent their respective counterparts.
With the hard constraint imposed in problem (11), the quantile regression curves fitted on finite-sample data do not cross. However, the inequality constraints introduced by this formulation may largely increase the computational cost. As such, we propose using a soft crossing penalty. This approach does not have any inequality constraint. Moreover, it provides practitioners with some flexibility to tolerate a certain level of crossing. Specifically, our NCKQR is defined as:
@@ 4.1 Kernel quantile regression
We compare our fastkqr algorithm with the R package kernlab (Karatzoglou et al., 2004) and the two generic optimizers nlm and optim in the R package stats. We chose the radial basis kernel and explored various combinations of the sample size $n$ and dimension $p$. We considered three different quantile levels: $\tau=0.1$, $0.5$, and $0.9$. For each training data, we applied the four solvers to fit KQR over 50 $\lambda$ values. The optimal tuning parameter $\lambda$ was selected using five-fold cross-validation, and the whole run time was recorded. The selected $\lambda$ was then used to compute the objective value of problem (2). The reported results represent averages from 20 independent repetitions. All computations were carried out on an Intel Core i5 (2.30 GHz) processor.
We use the same simulation model in Friedman et al. (2010). Predictors are generated from $\mathrm{N}(0,1)$, where each pair is correlated with $\rho=0.1$. The response values were generated by
$Y=\sum_{j=1}^{p}X_{j}\beta_{j}+cZ,$ (20)
where $\beta_{j}=(-1)^{j}\exp(-\frac{j-1}{10})$, $Z\sim\mathrm{N}(0,1)$, and $c$ is set so that the signal-to-noise ratio is 3.0.
$\tau$ $n$ fastkqr kernlab nlm optim
0.1 200 obj $0.553_{(0.091)}$ $0.553_{(0.091)}$ $0.555_{(0.087)}$ $0.598_{(0.048)}$
time 3.06 98.01 444.90 885.11
500 obj $0.591_{(0.048)}$ $0.591_{(0.048)}$ $0.592_{(0.048)}$ $0.611_{(0.044)}$
time 24.75 483.48 6669.63 13798.18
1000 obj $0.294_{(0.052)}$ $0.294_{(0.052)}$ $0.294_{(0.051)}$ $0.332_{(0.061)}$
time 114.97 2168.42 42013.01 $>24$h
0.5 200 obj $1.380_{(0.116)}$ $1.380_{(0.116)}$ $1.381_{(0.115)}$ $1.400_{(0.097)}$
time 2.63 106.75 423.24 858.95
500 obj $0.687_{(0.155)}$ $0.687_{(0.155)}$ $0.690_{(0.146)}$ $0.771_{(0.177)}$
time 24.44 593.39 6096.10 12350.00
1000 obj $0.868_{(0.270)}$ $0.868_{(0.270)}$ $0.869_{(0.270)}$ $0.954_{(0.301)}$
time 111.35 2382.14 43142.23 $>24$h
0.9 200 obj $0.522_{(0.119)}$ $0.522_{(0.119)}$ $0.527_{(0.111)}$ $0.577_{(0.076)}$
time 2.78 90.69 556.54 1124.95
500 obj $0.260_{(0.082)}$ $0.260_{(0.082)}$ $0.267_{(0.073)}$ $0.314_{(0.089)}$
time 23.33 445.50 6610.01 13333.11
1000 obj $0.304_{(0.051)}$ $0.304_{(0.051)}$ $0.306_{(0.048)}$ $0.398_{(0.202)}$
time 124.36 2282.20 43501.72 $>24$h
@@
Table 1: Objective values and computation time of four kernel quantile regression solvers for simulation data with $p=5000$, $n=\{200,500,1000\}$, and $\tau=\{0.1,0.5,0.9\}$. The numbers are the average quantities over 20 independent runs and the standard errors are presented in the parentheses.
@@
Tables 1 displays the average objective values and computation time for the four solvers, with all the run time including both model training and tuning parameter selection. Our algorithm fastkqr consistently outperforms the other solvers in speed across all the examples. Specifically, with a sample size $n=200$ and dimension $p=5000$, fastkqr outpaces kernlab by a factor of thirty. We also note that fastkqr and kernlab roughly have the same objective values, which are significantly lower than the other two generic optimizers.
@@ 4.2 Non-crossing kernel quantile regression
In this section, we compare fastkqr with the R package cvxr (Fu et al., 2020), and nlm and optim for fitting NCKQR. In each scenario, we simultaneously fit three quantile curves with $\tau=0.1,0.5$, and $0.9$ using the four solvers. With 20 independent runs, we calculate the average objective values of problem (12) and the total run time across 50 $\lambda_{2}$ values.
$p$ $n$ fastkqr cvxr nlm optim
100 200 obj $1.307_{(0.132)}$ $1.361_{(0.261)}$ $1.818_{(1.537)}$ $3.497_{(0.568)}$
time 4.04 7138.04 6573.51 13071.21
500 obj $2.529_{(0.127)}$ $2.612_{(0.331)}$
time 31.88 12739.79 $>24$h $>24$h
1000 obj $3.380_{(0.114)}$
time 222.12 $>24$h $>24$h $>24$h
1000 200 obj $1.121_{(0.102)}$ $1.166_{(0.225)}$ $3.219_{(1.414)}$ $4.277_{(0.567)}$
time 2.48 7067.55 6464.77 13036.94
500 obj $2.343_{(0.119)}$ $2.409_{(0.287)}$
time 28.61 15596.45 $>24$h $>24$h
1000 obj $3.372_{(0.132)}$
time 138.58 $>24$h $>24$h $>24$h
5000 200 obj $1.118_{(0.121)}$ $1.163_{(0.257)}$ $3.145_{(1.512)}$ $4.322_{(0.702)}$
time 3.24 7098.27 6328.77 12838.61
500 obj $2.290_{(0.134)}$ $2.352_{(0.278)}$
time 19.56 13180.66 $>24$h $>24$h
1000 obj $3.375_{(0.114)}$
time 117.09 $>24$h $>24$h $>24$h
Table 2: Objective values and computation time of four NCKQR solvers for simulation data with $p=\{100,1000,5000\}$, $n=\{200,500,1000\}$, and $\tau=\{0.1,0.5,0.9\}$. The numbers are the average quantities over 20 independent runs, and the standard errors are presented in the parentheses. A result shown as a star “*" means the corresponding solver cannot output a solution due to some numerical issue.
Table 2 presents the average objective values and computation time. Our algorithm is the fastest and the most accurate. For example, when $n=200$, our algorithm took only four seconds, in contrast to cvxr which spent about two hours. Furthermore, when $n$ is increased to 1000, our algorithm spent just 200 seconds, while the other three competitors took over 24 hours.
@@ 5 Discussions
Appendix A Algorithm
@@
1:$\mathbf{y}$, $\mathbf{K}$, $\tau$, and $\lambda^{[1]}>\lambda^{[2]}>\ldots>\lambda^{[L]}$.
@@
1:$\mathbf{y}$, $\mathbf{K}$, $\tau_{1}<\tau_{2}<\ldots<\tau_{T}$, $\lambda_{1}^{[1]}>\lambda_{1}^{[2]}>\ldots>\lambda_{1}^{[L_{1}]}$, and $\lambda_{2}^{[1]}>\lambda_{2}^{[2]}>\ldots>\lambda_{2}^{[L_{2}]}$.
@@
18:               Calculate the update formulae (LABEL:eq:nckqr-formula1), (24) and (LABEL:eq:nckqr-formula3) from right to left.
@@
Appendix B The fast spectral technique for NCKQR
@@
    $\boldsymbol{\Sigma}_{\gamma,\lambda_{1},\lambda_{2}}^{-1}\left(\begin{array}[]{c}\mathbf{1}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{1}^{\top}\mathbf{q}^{(k)}_{\tau_{t}}\\ \mathbf{K}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}-n\lambda_{2}\mathbf{K}^{\top}\boldsymbol{\alpha}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{K}^{\top}\mathbf{q}^{(k)}_{\tau_{t}}\end{array}\right)$ (21)
@@
  • Case 2. When $12\leq t\leq T-$ 1, we have
    $\boldsymbol{\Sigma}_{\gamma,\lambda_{1},\lambda_{2}}^{-1}\left(\begin{array}[]{c}\mathbf{1}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{1}^{\top}\mathbf{q}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{1}^{\top}\mathbf{q}^{(k)}_{\tau_{t-1}}\\ \mathbf{K}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}-n\lambda_{2}\mathbf{K}^{\top}\boldsymbol{\alpha}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{K}^{\top}\mathbf{q}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{K}^{\top}\mathbf{q}^{(k)}_{\tau_{t-1}}\end{array}\right)$ (24)
@@
    $\quad\cdot\left(\begin{array}[]{c}1\\ -\mathbf{v}\end{array}\right)+\left(\begin{array}[]{c}0\\ \mathbf{U}\boldsymbol{\Pi}_{\gamma,\lambda_{1},\lambda_{2}}^{-1}\boldsymbol{\Lambda}\mathbf{U}^{\top}\left(\mathbf{z}^{(k)}_{\tau_{t}}+n\lambda_{2}\boldsymbol{\alpha}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{q}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{q}^{(k)}_{\tau_{t-1}}\right)\end{array}\right).$ (29)
@@
    $\boldsymbol{\Sigma}_{\gamma,\lambda_{1},\lambda_{2}}^{-1}\left(\begin{array}[]{c}\mathbf{1}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{1}^{\top}\mathbf{q}^{(k)}_{\tau_{t-1}}\\ \mathbf{K}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}-n\lambda_{2}\mathbf{K}^{\top}\boldsymbol{\alpha}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{K}^{\top}\mathbf{q}^{(k)}_{\tau_{t-1}}\end{array}\right)$ (30)
@@
Appendix C Additional Simulations
@@
Table 3 presents the objective values and computation times using simulation data for $p=100$. Notably, fastkqr consistently emerges as the fastest solver, outperforming the other three solvers by at least an order of magnitude, while also achieving the lowest objective value.
$\tau$ $n$ fastkqr kernlab nlm optim
0.1 200 obj $0.551_{(0.077)}$ $0.551_{(0.077)}$ $0.558_{(0.066)}$ $0.583_{(0.051)}$
time 0.85 13.09 376.21 788.85
500 obj $0.553_{(0.045)}$ $0.553_{(0.045)}$ $0.566_{(0.039)}$ $0.588_{(0.037)}$
time 5.74 120.46 5351.71 10447.53
1000 obj $0.558_{(0.025)}$ $0.558_{(0.025)}$ $0.576_{(0.022)}$ $0.591_{(0.024)}$
time 20.91 695.72 45076.91 $>24$h
0.5 200 obj $1.292_{(0.093)}$ $1.292_{(0.093)}$ $1.297_{(0.089)}$ $1.332_{(0.081)}$
time 0.80 12.17 319.15 686.34
500 obj $1.287_{(0.054)}$ $1.287_{(0.054)}$ $1.301_{(0.050)}$ $1.331_{(0.049)}$
time 6.50 153.95 5356.08 10480.23
1000 obj $1.305_{(0.033)}$ $1.305_{(0.033)}$ $1.320_{(0.031)}$ $1.344_{(0.034)}$
time 17.97 686.88 46135.44 $>24$h
0.9 200 obj $0.541_{(0.072)}$ $0.541_{(0.072)}$ $0.547_{(0.065)}$ $0.579_{(0.054)}$
time 0.83 11.67 383.48 798.80
500 obj $0.541_{(0.050)}$ $0.541_{(0.050)}$ $0.557_{(0.040)}$ $0.584_{(0.034)}$
time 5.04 103.57 5156.50 10968.19
1000 obj $0.559_{(0.028)}$ $0.559_{(0.028)}$ $0.576_{(0.023)}$ $0.591_{(0.026)}$
time 22.14 709.62 44067.48 $>24$h

### v2

@@ 3.1 Methodology
When KQR is fitted at multiple quantile levels individually, say, for $0<\tau_{1}<\tau_{2}<\cdots<\tau_{T}<1$,
@@
the fitted curves may cross each other. To avoid the occurrence of crossing, Takeuchi et al. (2006)Bondell et al. (2010), and Liu and Wu (2011) consider fitting all the quantile curves simultaneously with a hard non-crossing constraint,
@@
where the notation ${\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}}_{t=1}^{T}$ represents the collection of $b_{\tau_{1}},\boldsymbol{\alpha}_{\tau_{1}},b_{\tau_{2}},\boldsymbol{\alpha}_{\tau_{2}},\ldots,b_{\tau_{T}},\boldsymbol{\alpha}_{\tau_{T}}$ for ease of presentation. In the sequel, we extend this notation to ${\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}}_{t=1}^{T}$ and ${\{\tilde{b}_{\tau_{t}},\tilde{\boldsymbol{\alpha}}_{\tau_{t}}\}}_{t=1}^{T}$ to represent their respective counterparts.
With the hard constraint imposed in problem (13), the quantile regression curves fitted on finite-sample data do not cross. However, the inequality constraints introduced by this formulation may largely increase the computational cost. As such, we propose using a soft crossing penalty. This approach does not have any inequality constraint. Moreover, it provides practitioners with some flexibility to tolerate a certain level of crossing. Specifically, our NCKQR is defined as
@@ 4.1 Kernel quantile regression
We compare our fastkqr algorithm with the Alternating Direction Method of Multipliers (ADMM) algorithm (Boyd et al., 2011), the optimizer kqr in the R package kernlab (Karatzoglou et al., 2004), the R packages clarabel (Goulart and Chen, 2024) and osqp (Stellato et al., 2020), and the two generic optimizers nlm and optim in the R package stats. We choose the radial basis kernel and explore various combinations of the sample size $n$ and dimension $p$. We consider three different quantile levels: $\tau=0.1$, 0.50.50.5, and 0.90.90.9. For each training data, we apply the seven solvers to fit KQR over 50 $\lambda$ values. The optimal tuning parameter $\lambda$ is selected using five-fold cross-validation, and the whole run time is recorded. The selected $\lambda$ is then used to compute the objective value of problem (2). The reported results represent averages from 20 independent repetitions. All computations are carried out on an Apple M1 (16GB) processor.
$\tau$ $n$ fastkqr kernlab clarabel ADMM osqp nlm optim
0.1 200 obj $0.365_{(0.044)}$ $0.365_{(0.044)}$ $0.365_{(0.044)}$ $0.365_{(0.044)}$ $0.365_{(0.044)}$ $0.367_{(0.041)}$ $0.384_{(0.033)}$
time 2.27 5.78 23.69 150.95 8.71 214.60 450.95
500 obj $0.356_{(0.019)}$ $0.356_{(0.019)}$ $0.356_{(0.019)}$ $0.356_{(0.019)}$ $0.356_{(0.019)}$ $0.358_{(0.019)}$ $0.368_{(0.020)}$
time 9.46 57.32 363.70 2479.34 149.68 3263.86 6671.74
1000 obj $0.360_{(0.012)}$ $0.360_{(0.012)}$ $0.360_{(0.012)}$ $0.360_{(0.012)}$ $0.360_{(0.012)}$ $0.361_{(0.012)}$ $0.366_{(0.011)}$
time 28.20 408.48 3802.69 10762.72 646.65 26906.18 54851.15
0.5 200 obj $0.813_{(0.063)}$ $0.813_{(0.063)}$ $0.813_{(0.063)}$ $0.813_{(0.063)}$ $0.813_{(0.063)}$ $0.822_{(0.057)}$ $0.844_{(0.053)}$
time 2.56 5.53 19.65 218.50 8.76 221.58 449.92
500 obj $0.807_{(0.044)}$ $0.807_{(0.044)}$ $0.807_{(0.044)}$ $0.807_{(0.044)}$ $0.807_{(0.044)}$ $0.813_{(0.041)}$ $0.827_{(0.040)}$
time 9.56 54.83 285.14 2654.29 141.39 3328.60 6804.10
1000 obj $0.811_{(0.024)}$ $0.811_{(0.024)}$ $0.811_{(0.024)}$ $0.811_{(0.024)}$ $0.812_{(0.024)}$ $0.816_{(0.022)}$ $0.829_{(0.020)}$
time 28.44 368.38 2711.26 10784.24 661.21 27568.66 55697.12
0.9 200 obj $0.380_{(0.041)}$ $0.380_{(0.041)}$ $0.380_{(0.041)}$ $0.380_{(0.041)}$ $0.380_{(0.041)}$ $0.384_{(0.038)}$ $0.403_{(0.034)}$
time 2.29 5.09 26.36 121.39 7.35 224.17 455.74
500 obj $0.377_{(0.032)}$ $0.377_{(0.032)}$ $0.377_{(0.032)}$ $0.377_{(0.032)}$ $0.377_{(0.032)}$ $0.378_{(0.031)}$ $0.394_{(0.037)}$
time 9.57 53.47 377.31 2344.44 125.50 3326.35 6709.23
1000 obj $0.365_{(0.010)}$ $0.365_{(0.010)}$ $0.365_{(0.010)}$ $0.365_{(0.010)}$ $0.365_{(0.010)}$ $0.366_{(0.010)}$ $0.403_{(0.046)}$
time 27.98 386.04 3785.58 10346.09 622.64 27715.45 55714.28
@@
Table 1: Objective values and computation time of seven kernel quantile regression solvers for simulation data (Yuan, 2006) with $p=2$, $n=\{200,500,1000\}$, and $\tau=\{0.1,0.5,0.9\}$. The numbers are the average quantities over 20 independent runs and the standard errors are presented in the parentheses.
$n$ fastkqr cvxr nlm optim
200 obj $2.527_{(0.108)}$ $2.630_{(0.105)}$ $3.048_{(0.105)}$ $3.127_{(0.135)}$
time 3.37 1149.93 2938.42 6392.16
500 obj $2.730_{(0.112)}$ $2.998_{(0.162)}$ $3.423_{(0.132)}$
time 17.06 3439.39 41266.57 >24>24> 24h
1000 obj $3.380_{(0.114)}$
time 57.56 20171.01 >24>24> 24h >24>24> 24h
Table 2: Objective values and computation time of four NCKQR solvers for simulation data (Yuan, 2006) with $p=2$, $n=\{200,500,1000\}$, and $\tau=\{0.1,0.5,0.9\}$. The average quantities over 20 independent runs are displayed, and standard errors are presented in parentheses. A result shown as a star “*” means the corresponding solver cannot output a solution due to numerical issues.
Following Yuan (2006), two-dimensional data are generated based on
where $X_{1}$ and $X_{2}$ are drawn from Uniform (0,1)(0,1)( 0 , 1 ), and the error term $\epsilon$ is from the standard normal distribution.
@@
Tables 1 displays the average objective values and computation time for the seven solvers, with all the run time including both model training and tuning parameter selection. Our algorithm fastkqr consistently outperforms the other solvers in speed across all the examples. For instance, when $n=1000$, our fastkqr algorithm was more than an order of magnitude faster than kernlab, and notably, more than 400 times faster than nlm and optim. As the sample size $n$ grows, the computational advantages of our algorithm over the others become more significant. Additionally, apart from nlm and optim, the other five solvers exhibit roughly the same objective values, which are notably lower than those of the two generic optimizers.
@@ 4.2 Non-crossing kernel quantile regression
In this section, we compare fastkqr with the R package cvxr (Fu et al., 2020), and nlm and optim for fitting NCKQR. In each scenario, we simultaneously fit three quantile curves with $\tau=0.1,0.5$, and 0.90.90.9 using the four solvers. With 20 independent runs, we calculate the average objective values of problem (14) and the total run time across 50 $\lambda_{2}$ values.
Table 2 presents the average objective values and computation time. Our algorithm is the fastest and the most accurate. For example, when $n=500$, our algorithm took only seventeen seconds, in contrast to cvxr which spent about one hour. Furthermore, when $n$ is increased to 1000, our algorithm spent just 57 seconds, while nlm and optim took over 24 hours.
@@ 5 Discussions
6 Acknowledgments
The authors extend their sincere gratitude to the Editor, the Associate Editor, and the two anonymous Referees for their insightful and constructive comments, which have greatly improved the quality of our work.
7 Disclosure Statement
The authors report there are no competing interests to declare.
@@ 7 Disclosure Statement
Appendix S1 Plot of $\gamma$-Smoothed Check Loss
Figure S1 illustrates the $\gamma$-smoothed check loss for various values of $\gamma$, along with the original check loss.
[…]
Appendix S2 Algorithms
@@
[…]
@@
[…]
@@
18:     Calculate the update formulae (LABEL:eq:nckqr-formula1), (S26) and (LABEL:eq:nckqr-formula3) from right to left.
@@
Appendix S3 The Fast Spectral Technique for NCKQR
@@
    $\boldsymbol{\Sigma}_{\gamma,\lambda_{1},\lambda_{2}}^{-1}\left(\begin{array}[]{c}\mathbf{1}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{1}^{\top}\mathbf{q}^{(k)}_{\tau_{t}}\\ \mathbf{K}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}-n\lambda_{2}\mathbf{K}^{\top}\boldsymbol{\alpha}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{K}^{\top}\mathbf{q}^{(k)}_{\tau_{t}}\end{array}\right)$ (S23)
@@
  • Case 2. When $2\leq t\leq T-1$, we have
    $\boldsymbol{\Sigma}_{\gamma,\lambda_{1},\lambda_{2}}^{-1}\left(\begin{array}[]{c}\mathbf{1}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{1}^{\top}\mathbf{q}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{1}^{\top}\mathbf{q}^{(k)}_{\tau_{t-1}}\\ \mathbf{K}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}-n\lambda_{2}\mathbf{K}^{\top}\boldsymbol{\alpha}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{K}^{\top}\mathbf{q}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{K}^{\top}\mathbf{q}^{(k)}_{\tau_{t-1}}\end{array}\right)$ (S26)
@@
    $\quad\cdot\left(\begin{array}[]{c}1\\ -\mathbf{v}\end{array}\right)+\left(\begin{array}[]{c}0\\ \mathbf{U}\boldsymbol{\Pi}_{\gamma,\lambda_{1},\lambda_{2}}^{-1}\boldsymbol{\Lambda}\mathbf{U}^{\top}\left(\mathbf{z}^{(k)}_{\tau_{t}}+n\lambda_{2}\boldsymbol{\alpha}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{q}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{q}^{(k)}_{\tau_{t-1}}\right)\end{array}\right).$ (S31)
@@
    $\boldsymbol{\Sigma}_{\gamma,\lambda_{1},\lambda_{2}}^{-1}\left(\begin{array}[]{c}\mathbf{1}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{1}^{\top}\mathbf{q}^{(k)}_{\tau_{t-1}}\\ \mathbf{K}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}-n\lambda_{2}\mathbf{K}^{\top}\boldsymbol{\alpha}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{K}^{\top}\mathbf{q}^{(k)}_{\tau_{t-1}}\end{array}\right)$ (S32)
@@
Appendix S4 Alternative smoothing surrogates
Moreau envelope. This approach smooths a function by introducing a quadratic regularization term. Given a function $f(t)$, the Moreau envelope $f_{\gamma}(t)$ is defined as $f_{\gamma}(t)=\inf_{y}\left\{f(y)+\frac{1}{2\gamma}\|t-y\|_{2}^{2}\right\}.$ Applying this concept to the check loss function $\rho_{\tau}(t)$ yields
$H_{\gamma,\tau}(t)=\inf_{y}\left\{\rho_{\tau}(y)+\frac{1}{2\gamma}\|t-y\|_{2}^{2}\right\}=\begin{cases}t(\tau-1)-\frac{1}{2}(\tau-1)^{2}\gamma&\text{ if }\quad t<(\tau-1)\gamma,\\ \frac{t^{2}}{2\gamma}&\text{ if }\quad(\tau-1)\gamma\leq t\leq\tau\gamma,\\ t\tau-\frac{1}{2}\tau^{2}\gamma&\text{ if }t>\tau\gamma.\end{cases}$
Nesterov’s smoothing.  Consider a function $f(t)$, Nesterov’s smoothing approach (Nesterov, 2005) constructs a smooth approximation $f_{\gamma}(t)=\sup_{x\in\operatorname{dom}(g)}\langle t,x\rangle-(g(x)+\gamma d(x)),$ where $g(x)$ is the conjugate function of $f(x)$ and $d(x)$ is a prox-function.
Applying this to the check loss $\rho_{\tau}(t)$ with $g(x)=\max_{z\in\mathbb{R}}\{xz-z(p-\mathbf{1}(z<0))\}$ and setting $d(x)=\frac{1}{2}\|x\|_{2}^{2}$ yields
$H_{\gamma,\tau}(t)=\left\{\begin{array}[]{ll}t(\tau-1)-\frac{1}{2}(\tau-1)^{2}\gamma&\text{if}\enskip{}t\leq(\tau-1)\gamma,\\ \frac{t^{2}}{2\gamma}&\text{if}\enskip{}(\tau-1)\gamma\leq{t}\leq\tau\gamma,\\ t\tau-\frac{1}{2}\tau^{2}\gamma&\text{if}\enskip{}t\geq\tau\gamma.\end{array}\right.$
[…]
which coincides exactly with the proposed smooth check loss function up to a constant shift.
[…]
  1. 1.
    (Uniform kernel) Consider the uniform kernel $K(u)=(1/2)I(|u|\leq 1)$, substituting this into the above integral yields
[…]
    which is exactly the proposed smoothed check loss function.
  2. 2.
    (Epanechnikov kernel) Consider the Epanechnikov kernel $K(u)=(3/4)\left(1-u^{2}\right)I(|u|\leq 1)$, then the smoothed check loss takes the form
[…]
It can be easily shown that all the aforementioned smoothed loss functions satisfy the gradient conditions stated in Theorem (5); as such, they can recover the exact solution through the finite smoothing algorithm.
Appendix S5 Additional Simulation Results
In this section, we use the same simulation model in Friedman et al. (2010) to further demonstrate the performance of fastkqr. Predictors are generated from $\mathrm{N}(0,1)$, where each pair is correlated with $\rho=0.1$. The response values were generated by
$Y=\sum_{j=1}^{p}X_{j}\beta_{j}+cZ,$ (S33)
where $\beta_{j}=(-1)^{j}\exp(-\frac{j-1}{10})$, $Z\sim\mathrm{N}(0,1)$, and $c$ is set so that the signal-to-noise ratio is 3.0. All computations were carried out on an Apple M1 (16GB) processor.
@@
Table S1 and Table S2 present the objective values and computation time using simulation data for $p=100$ and 500050005000, respectively. Notably, fastkqr consistently emerges as the fastest solver, outperforming the other three solvers by at least an order of magnitude, while also achieving the lowest objective value.
$\tau$ $n$ fastkqr kernlab clarabel ADMM osqp nlm optim
0.1 200 obj $0.601_{(0.054)}$ $0.601_{(0.054)}$ $0.601_{(0.054)}$ $0.601_{(0.054)}$ $0.601_{(0.054)}$ $0.601_{(0.054)}$ $0.611_{(0.051)}$
time 0.31 5.87 21.43 174.99 8.92 217.40 452.91
500 obj $0.563_{(0.033)}$ $0.563_{(0.033)}$ $0.563_{(0.033)}$ $0.563_{(0.033)}$ $0.563_{(0.033)}$ $0.573_{(0.031)}$ $0.593_{(0.032)}$
time 1.87 53.58 328.30 2368.89 162.54 3492.46 6785.29
1000 obj $0.539_{(0.019)}$ $0.539_{(0.019)}$ $0.539_{(0.019)}$ $0.539_{(0.019)}$ $0.539_{(0.019)}$ $0.561_{(0.016)}$ $0.578_{(0.024)}$
time 9.07 368.41 3264.13 10643.25 635.71 28046.38 56038.72
0.5 200 obj $0.939_{(0.224)}$ $0.939_{(0.224)}$ $0.939_{(0.224)}$ $0.939_{(0.224)}$ $0.939_{(0.224)}$ $1.021_{(0.162)}$ $1.157_{(0.131)}$
time 0.28 6.40 17.07 294.33 10.43 212.05 450.62
500 obj $0.953_{(0.144)}$ $0.953_{(0.144)}$ $0.953_{(0.144)}$ $0.953_{(0.144)}$ $0.953_{(0.144)}$ $1.089_{(0.100)}$ $1.218_{(0.075)}$
time 1.62 55.06 243.30 2441.78 175.92 3215.64 6732.79
1000 obj $1.009_{(0.087)}$ $1.009_{(0.087)}$ $1.009_{(0.087)}$ $1.009_{(0.087)}$ $1.009_{(0.087)}$ $1.154_{(0.046)}$ $1.232_{(0.055)}$
time 7.93 362.85 2569.78 10827.59 770.64 27356.19 54737.99
0.9 200 obj $0.585_{(0.041)}$ $0.585_{(0.041)}$ $0.585_{(0.041)}$ $0.585_{(0.041)}$ $0.585_{(0.041)}$ $0.586_{(0.040)}$ $0.598_{(0.038)}$
time 0.32 5.54 21.90 162.68 9.14 219.43 451.78
500 obj $0.554_{(0.029)}$ $0.554_{(0.029)}$ $0.554_{(0.029)}$ $0.554_{(0.029)}$ $0.554_{(0.029)}$ $0.566_{(0.026)}$ $0.591_{(0.037)}$
time 1.87 51.74 324.96 2424.24 161.57 3363.38 6746.65
1000 obj $0.541_{(0.017)}$ $0.541_{(0.017)}$ $0.541_{(0.017)}$ $0.541_{(0.017)}$ $0.541_{(0.017)}$ $0.565_{(0.016)}$ $0.581_{(0.016)}$
time 8.89 366.14 3599.66 10632.72 620.84 28308.27 55795.43